OpenStudy (anonymous):

Ho Do You Write 4/5 As A Percent ? @dmezzullo

thomaster (thomaster):

when 5 = 100%, what is 4 in %?

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

multiply by 20 so (4/5)20 = 80/100 so answer is?

OpenStudy (ryan123345):

@Luis_Rivera how can you say 4/5 is 40% when 2.5/5= 50% ? lol

OpenStudy (anonymous):

@Luis_Rivera what is 4*20?

OpenStudy (ryan123345):

since when did 4*20 = 40? lol @Luis_Rivera

terenzreignz (terenzreignz):